OpenStudy (anonymous):

A snare drum has diameter fourteen inches. A bass drum has diameter thirty-six inches. What is the difference in the areas of the two drums? sixty-nine point zero eight square inches four hundred eighty-four square inches eight hundred sixty-three point five square inches three thousand nine hundred fifteen point five eight square inches

OpenStudy (campbell_st):

ok... so the area of a circle is \[A = \pi r^2\] snare has a radius or 7 inches bass a radius of 18 inches to the difference in area is \[Diff = \pi \times 18^2 - \pi \times 7^2\]

OpenStudy (anonymous):

Ok so which one is it?